the sum of the ages of a brother and sister is 28. if four time the brothers age is subtracted from three times the sister age , the difference is 14
Answer by Timnewman(323)

Let the age of the brother and sister be x and y.
from your question,
x+y=28----(1)
3x-4y=14----(2)
from 1,
x=28-y---(3)
put 3 in 2
3(28-y)-4y=14
84-3y-4y=14
7y=70
y=10
put y=10 in 3
x=28-10
x=18
The ages of sister and brother is 18yrs and 10yrs respectively.
HOPE THIS HELPS?